OpenStudy (anonymous):

if I'm finding the inverse of f(x) = 6 - 5x^2, do I take the square root of -5x^2 or 5x^2?

OpenStudy (jdoe0001):

do the same as before hmmm first, swap the variables

OpenStudy (anonymous):

right, did that. just wondering what side of the minus I put the square root on. cuz I know that changes things

OpenStudy (jdoe0001):

well, what you can do is, multiply both sides by -1, to make that positive or just add \(5y^2\) to both sides but notice, the 5 is not part of the square

OpenStudy (jdoe0001):

\(\bf f(x)={\color{brown}{ y}}=6-5{\color{blue}{ x}}^2\qquad \qquad inverse\implies {\color{blue}{ x}}=6-5{\color{brown}{ y}}^2 \\ \quad \\ x=6-5y^2\implies x=6-5\cdot y^2\implies x-6=-5\cdot y^2 \\ \quad \\ \cfrac{x-6}{-5}=y^2\implies \sqrt{\cfrac{x-6}{-5}}=y\)

OpenStudy (anonymous):

oh how not? n I'll just multiply them by -1